# Mitigating Omitted Variable Bias in Empirical Software Engineering

### v3

@@ 4.4.1 Simulation Parameters
Instead, we go back to Feldt et al. [18] and use the statistics from some of the reviewed primary studies to get a ballpark estimate of the relevant effects. This trades off some generality for a manageable simulation time. Table 4 shows the range of parameter values that we used for our simulations.
@@
First of all, we want to simulate all plausible confounding scenarios of $Z$; therefore, we consider all combinations of small (0.1), medium (0.3), and large (0.5) standardized effect sizes [^16] The values 0.1, 0.3, and 0.5 are often considered as the boundaries between negligible, small, medium, and large effect sizes for effects measured as standardized mean differences [13, p. 224–225] [19]. This is just a plausible choice of values for our illustrative example; in concrete case studies, one could pick parameter values more precisely, based on what is considered a small, medium, or large effect in the domain of interest. —both positive and negative—for each parameter $\gamma_{e}$ and $\gamma_{t}$.
@@
We also consider all possible effect sizes for parameter $t_{e}$, which represents the treatment effect that we are trying to estimate; however, we only consider positive effect sizes since, according to the studies reviewed in Feldt et al. [18], it is implausible that large teams produce an overall lesser effort than small teams.
@@
As for the effects $x_{e}$ corresponding to the edges $X\to E$, for every other variable $X$ in the adjustment set, we picked a small, medium, or large, positive or negative effect size based on some of the studies reviewed in Feldt et al. [18]. Since those studies focused on effort (or related variables) as outcome, we could not find any hard data about the magnitude of the effects of these other variables $X$ on other covariates. Simplistically, we assume that $x_{y}$ is the same as $x_{e}$, that is variable $X$ has roughly the same effect on all variables it directly affects. Again, a specific case study could come up with more definite estimates; our goal is mainly to demonstrate this analysis method on somewhat plausible data.
@@
$k_{e}$ $K\to E$ 0.1 [78, Tab. XIII, % variance explained/project type]
@@ 4.4.2 Simulation Process
This sample/fit process is repeated $n_{\textrm{sim}}$ times for each parameter combination; finally, Algorithm 1 returns the mean $\beta_{t}$, the 50% highest-probability density interval $\ell_{50}\ldots u_{50}$, and the 95% highest-probability density interval $\ell_{95}\ldots u_{95}$ of the estimates of $\beta_{t}$ over all repetitions. [^17] Highest density intervals are a Bayesian analog to confidence intervals in frequentist statistics [24, 40]. Since highest density intervals are simply defined as intervals on a probability distribution, their interpretation is usually considered more intuitive [33, 30]. These statistics summarize the likely ranges of $\beta_{t}$ estimated from the simulations.
@@ 4.4.3 Simulation Results
Figures 8, 9, and 10 display (a significant subset of) the simulation results. Each figure includes 16 $4\times 4$ subplots for each combination of the following values for $\gamma_{e}$ and $\gamma_{t}$: -0.5, 0.1, 0.3, 0.5. Figure 8 refers to the experiments where the effect $t_{e}$ to be estimated is small (0.1); Figure 9 where $t_{e}$ is medium (0.3); and Figure 9 where $t_{e}$ is large (0.5). Overall, the subset of experimental results displayed in the three figures is sufficient to see the main trends in the sensitivity analysis; anyway, the replication package [23] includes plots for all parameter combinations. The capsule summary of these results is that it is possible to retrieve a reasonably precise estimate of the effect $t_{e}$ provided the confounding effect of the unmeasured $Z$ (parameters $\gamma_{e}$ and $\gamma_{t}$) is not large compared to $t_{e}$.
@@ 4.5 Sensitivity Analysis with E-values
The sensitivity analysis based on simulation presented in Section 4.4 is informative and flexible, but it also has clear disadvantages: it can be very time consuming, and it does usually requires a good amount of empirical data to tune the simulation parameters in a way that is representative of the domain. As a less demanding alternative, this section presents a tipping-point analysis similar to the one we described in Section 3.2 for the other illustrative example.
@@ 4.5.1 E-values
Section 3.2’s sensitivity analysis uses the scd (scaled-mean difference) as a standardized measure of the effect of an unmeasured confounder on the treatment. The scd (1) is defined based on a dichotomous partition of the treatment variable. In Section 3’s domain, the treatment was the programming language, which is naturally dichotomous. In contrast, team size $T$ (the treatment variable in our current domain) is an intrinsically quantitative (numeric) variable; in order to calculate an scd of $Z\to T$, we would have to arbitrarily partition teams into small vs. large. Instead, we rely on a different kind of tipping-point analysis based on the notion of E-value [75], which is also applicable to continuous quantitative exposure variables. [^18] We rely on the R package EValue [49], which implements a variety of state-of-the-art sensitivity analysis techniques for unmeasured confounding based on the notion of E-value [11, 75, 74].
The E-value [^19] The “E” stands for “Evidence”. is “the minimum strength of association, on the risk ratio scale, that an unmeasured confounder would need to have with both the treatment and the outcome to fully explain away a specific treatment-outcome association, conditional on the measured covariates” [75]. In our scenario, an estimate $\beta_{t}$ of the observed effect $T\to E$, obtained by fitting model (4) on empirical data, would represent the “specific treatment-outcome association, conditional on the measured covariates.” The E-value can be computed from $\beta_{t}$, as well as $\epsilon_{t}$ (the standard error of $\beta_{t}$’s estimate), an estimate of $\sigma$ (parameter $\sigma$ in model (4)), and a parameter $\delta$ that quantifies the arbitrarily chosen change in treatment variable $T$. With these parameters, an E-value of $e$ can be interpreted as follows: if the unmeasured confounder $Z$ were strong enough to increase, by a factor of $e$, the probability of raising the exposure $T$ by $\delta$—while simultaneously affecting the outcome $E$ by a similar amount in standardized units—then the observed effect $\beta_{t}$ would be entirely due to $Z$’s confounding.
@@ 4.5.2 Computing E-values
Let’s get into computing the E-value in our illustrative example of team size and effort. Since we don’t have real-world data directly about the variables of Figure 7, we resort, once again, to simulation to get some plausible data in a convenient format. Unlike in Section 4.4, now we don’t need to perform many repetitions with small sample size, since the simulated data will not be used for a sensitivity analysis but to create a synthetic dataset based on the Table 4’s parameters. Thus, we simply draw ten thousand datapoints by sampling model (3) for each of the following parameter combinations:
@@ 4.5.3 Confounding Sensitivity
    When $t_{e}=0.5$ and $\delta=0.1$, the E-value is $1.26$. Informally, to explain away a large observed effect, $Z$ should increase by 26% the probability of introducing a small increase in team size ($\delta=0.1$). This is definitely a more significant impact than in the previous point i), but it is perhaps still plausible.
@@
    When $t_{e}=\delta=0.5$, the E-value is $1.80$. Informally, to explain away a large observed effect, $Z$ should increase by 80% the probability of introducing a large increase in team size. This scenario is no longer so plausible: it is a case of “extraordinary claims [requiring] extraordinary evidence” [64].
Comparing scenarios ii) and iii) is also interesting. Overall, they suggest that it is still possible that a confounder is responsible for a large effect; however, this is plausible only if the confounder $Z$ can bias the estimate of the $T\to E$ effect with only a small change in the treatment $T$ (scenario ii), where the $T\to E$ is more sensitive), as opposed to a large change (which is less likely to have been missed by the quantitative analysis).
As usual, with a better understanding of the domain (for example, the typical project characteristics in the company where the analyzed data was collected), the E-value analysis could produce more actionable results. For example, if one is confident that the estimates of team size and project effort that are normally produced are usually precise, it would indicate that it’s unlikely that a confounder would go undetected even for small changes of treatment/outcome (e.g., $\delta=0.1)$. Conversely, if the characteristics of the projects that are being analyzed make estimates intrinsically imprecise or uncertain, an unmeasured confounder that substantially increases the probability of swaying such estimates is a plausible possibility. As one example, Morasca and Russo [56] study of productivity (one of those surveyed in Feldt et al. [18]) presents estimates of productivity with a large dispersion ($\sigma\in[0.35,7.2]$); these indicate much uncertainty, thus raising the possibility of unmeasured confounders.
@@ 5 Dealing with Omitted Variable Bias
What kind of information we can use to build a DAG depends on the domain we are investigating, and on the maturity of the state of the art. If there are plenty of rigorous primary studies about the quantities of interest, and perhaps even systematic reviews or meta-studies, we can summarize their evidence in a DAG—similarly to what is proposed by Feldt et al. [18]. If our study’s target is novel or less established, we may have to rely on domain expertise and intuition to build a DAG. Even if there is a lot of uncertainty about the precise causal structure underlying a certain domain, there are techniques to (partially) validate candidate DAGs [22]. It is quite natural to also consider different possible DAGs, and to use them to perform a “what if” analysis in different scenarios. In this step of the analysis, causal DAGs are mainly a convenient notation to rigorously express our knowledge or hypotheses about the causal relations among variables, and to investigate their consequences on the overall results of our analysis.
@@
As shown in Section 4.2, given a DAG and treatment/outcome variables, one can systematically compute an adjustment set: a set of covariates that should be conditioned on in a regression to ensure that the coefficient associated with the treatment variable estimates the unbiased effect of the treatment on the outcome (“controlling for” the spurious influence of any confounders).
In the best-case scenarios, one can simply use an adjustment set to prevent confounding. Unfortunately, this is not always possible. First, the adjustment set’s validity is predicated on the accuracy and completeness of the DAG: if we missed some relevant variables, or misrepresented some causal relations, an adjustment set no longer guarantees an unbiased estimate. Second, even if we are confident the DAG is accurate, certain DAGs do not admit adjustment sets (because different kinds of confounding require incompatible adjustment sets [50]), or some variables in the adjustment set are hard or impossible to measure (a scenario that we explored in Section 4.3). In these cases, the next steps in this list can help deal with these shortcomings.
@@
Using the parameter estimates identified in the previous step, one can perform different kinds of sensitivity analyses of unmeasured confounder. These analyses provide a quantitative estimate (usually on a normalized scale) of how much some unmeasured confounder may bias the estimate of an effect of interest. In the paper, we demonstrated two kinds of so-called tipping-point sensitivity analysis, which express how strong an unmeasured confounder should be to cancel out an observed treatment/outcome effect. Section 3.2 presented a tipping-point analysis based on an estimate of the scaled-mean difference of an unmeasured confounder on the treatment, which is applicable to dichotomous treatment variables. Section 4.5 presented a sensitivity analysis based on E-values—a probability ratio between the confounded and non-confounded scenarios—which is also applicable to continuous treatment variables.
@@
If the previous step’s sensitivity analysis is inconclusive, in that it failed to rule out the possibility of confounding, one can perform a more precise analysis of confounding based on simulation. Section 4.4 illustrated this on our second illustrative example of teamwork productivity. A simulation analysis is flexible, because one can explore many different variants of generative and inference models. It also supports analyzing the impact of dealing with small sample sizes in a way that realistically reflects the availability of data in the study domain. These advantages come with a cost in terms of simulation time; usually, however, it still is much cheaper to perform a detailed simulation than to embark “blind” in running an empirical study without a clear understanding of the possible confounders, and of the threats to validity they may introduce.
@@ 6 Conclusions
This paper introduced to the empirical software engineering community techniques to assess and mitigate the so-called omitted variable bias. These techniques are crucially based on a causal model of the relations among variables of interest, formalized by means the causal DAG notation [58].
First, if the causal structure among variables admits an adjustment set that only includes measurable variables, one can correct for an omitted variable bias by simply conditioning on all variables in the adjustment set. If this is not possible, one can perform a sensitivity analysis, whose goal is investigating the impact of unknown confounders. The paper presented different kinds of sensitivity analyses, including both so-called tipping-point analyses based on canonical distributional assumptions, and more precise, but also computationally expensive, analyses based on simulation. We demonstrated these techniques on two illustrative examples—the relation between programming languages and code quality, and the effect on team size on software development effort—taken from recent statistical analyses of data in these two domains [60, 18].

### v4

@@ 4.4.1 Simulation Parameters
Instead, we go back to Feldt et al. (2025) and use the statistics from some of the reviewed primary studies to get a ballpark estimate of the relevant effects. This trades off some generality for a manageable simulation time. Table 4 shows the range of parameter values that we used for our simulations.
@@
First of all, we want to simulate all plausible confounding scenarios of $Z$; therefore, we consider all combinations of small (0.1), medium (0.3), and large (0.5) standardized effect sizes [^16] The values 0.1, 0.3, and 0.5 are often considered as the boundaries between negligible, small, medium, and large effect sizes for effects measured as standardized mean differences (Cohen, 1988, p. 224–225) Fey et al. (2023). This is just a plausible choice of values for our illustrative example; in concrete case studies, one could pick parameter values more precisely, based on what is considered a small, medium, or large effect in the domain of interest. —both positive and negative—for each parameter $\gamma_{e}$ and $\gamma_{t}$.
@@
We also consider all possible effect sizes for parameter $t_{e}$, which represents the treatment effect that we are trying to estimate; however, we only consider positive effect sizes since, according to the studies reviewed in Feldt et al. (2025), it is implausible that large teams produce an overall lesser effort than small teams.
@@
As for the effects $x_{e}$ corresponding to the edges $X\to E$, for every other variable $X$ in the adjustment set, we picked a small, medium, or large, positive or negative effect size based on some of the studies reviewed in Feldt et al. (2025). Since those studies focused on effort (or related variables) as outcome, we could not find any hard data about the magnitude of the effects of these other variables $X$ on other covariates. Simplistically, we assume that $x_{y}$ is the same as $x_{e}$, that is variable $X$ has roughly the same effect on all variables it directly affects. Again, a specific case study could come up with more definite estimates; our goal is mainly to demonstrate this analysis method on somewhat plausible data.
@@
$k_{e}$ $K\to E$ 0.1 (Wang et al., 2008, Tab. XIII, % variance explained/project type)
@@ 4.4.2 Simulation Process
This sample/fit process is repeated $n_{\textrm{sim}}$ times for each parameter combination; finally, Algorithm 1 returns the mean $\beta_{t}$, the 50% highest-probability density interval $\ell_{50}\ldots u_{50}$, and the 95% highest-probability density interval $\ell_{95}\ldots u_{95}$ of the estimates of $\beta_{t}$ over all repetitions. [^17] Highest density intervals are a Bayesian analog to confidence intervals in frequentist statistics Gelman et al. (2014); Jaynes (1976). Since highest density intervals are simply defined as intervals on a probability distribution, their interpretation is usually considered more intuitive Hoekstra et al. (2014); Hespanhol et al. (2019). These statistics summarize the likely ranges of $\beta_{t}$ estimated from the simulations.
@@ 4.4.3 Simulation Results
Figures 8, 9, and 10 display (a significant subset of) the simulation results. Each figure includes 16 $4\times 4$ subplots for each combination of the following values for $\gamma_{e}$ and $\gamma_{t}$: -0.5, 0.1, 0.3, 0.5. Figure 8 refers to the experiments where the effect $t_{e}$ to be estimated is small (0.1); Figure 9 where $t_{e}$ is medium (0.3); and Figure 9 where $t_{e}$ is large (0.5). Overall, the subset of experimental results displayed in the three figures is sufficient to see the main trends in the sensitivity analysis; anyway, the replication package Furia and Torkar (2025) includes plots for all parameter combinations. The capsule summary of these results is that it is possible to retrieve a reasonably precise estimate of the effect $t_{e}$ provided the confounding effect of the unmeasured $Z$ (parameters $\gamma_{e}$ and $\gamma_{t}$) is not large compared to $t_{e}$.
@@ 4.5 Sensitivity Analysis with E-values
The sensitivity analysis based on simulation presented in subsection 4.4 is informative and flexible, but it also has clear disadvantages: it can be very time consuming, and it does usually requires a good amount of empirical data to tune the simulation parameters in a way that is representative of the domain. As a less demanding alternative, this section presents a tipping-point analysis similar to the one we described in subsection 3.2 for the other illustrative example.
@@ 4.5.1 E-values
subsection 3.2’s sensitivity analysis uses the scd (scaled-mean difference) as a standardized measure of the effect of an unmeasured confounder on the treatment. The scd (1) is defined based on a dichotomous partition of the treatment variable. In section 3’s domain, the treatment was the programming language, which is naturally dichotomous. In contrast, team size $T$ (the treatment variable in our current domain) is an intrinsically quantitative (numeric) variable; in order to calculate an scd of $Z\to T$, we would have to arbitrarily partition teams into small vs. large. Instead, we rely on a different kind of tipping-point analysis based on the notion of E-value (VanderWeele and Ding, 2011), which is also applicable to continuous quantitative exposure variables. [^18] We rely on the R package EValue (Mathur et al., 2021), which implements a variety of state-of-the-art sensitivity analysis techniques for unmeasured confounding based on the notion of E-value (Chinn, 2000; VanderWeele and Ding, 2011; VanderWeele, 2017).
The E-value [^19] The “E” stands for “Evidence”. is “the minimum strength of association, on the risk ratio scale, that an unmeasured confounder would need to have with both the treatment and the outcome to fully explain away a specific treatment-outcome association, conditional on the measured covariates” VanderWeele and Ding (2011). In our scenario, an estimate $\beta_{t}$ of the observed effect $T\to E$, obtained by fitting model (4) on empirical data, would represent the “specific treatment-outcome association, conditional on the measured covariates.” The E-value can be computed from $\beta_{t}$, as well as $\epsilon_{t}$ (the standard error of $\beta_{t}$’s estimate), an estimate of $\sigma$ (parameter $\sigma$ in model (4)), and a parameter $\delta$ that quantifies the arbitrarily chosen change in treatment variable $T$. With these parameters, an E-value of $e$ can be interpreted as follows: if the unmeasured confounder $Z$ were strong enough to increase, by a factor of $e$, the probability of raising the exposure $T$ by $\delta$—while simultaneously affecting the outcome $E$ by a similar amount in standardized units—then the observed effect $\beta_{t}$ would be entirely due to $Z$’s confounding.
@@ 4.5.2 Computing E-values
Let’s get into computing the E-value in our illustrative example of team size and effort. Since we don’t have real-world data directly about the variables of Figure 7, we resort, once again, to simulation to get some plausible data in a convenient format. Unlike in subsection 4.4, now we don’t need to perform many repetitions with small sample size, since the simulated data will not be used for a sensitivity analysis but to create a synthetic dataset based on the Table 4’s parameters. Thus, we simply draw ten thousand datapoints by sampling model (3) for each of the following parameter combinations:
@@ 4.5.3 Confounding Sensitivity
    When $t_{e}=0.5$ and $\delta=0.1$, the E-value is $1.26$. Informally, to explain away a large observed effect, $Z$ should increase by 26% the probability of introducing a small increase in team size ($\delta=0.1$). This is definitely a more significant impact than in the previous point item i), but it is perhaps still plausible.
@@
    When $t_{e}=\delta=0.5$, the E-value is $1.80$. Informally, to explain away a large observed effect, $Z$ should increase by 80% the probability of introducing a large increase in team size. This scenario is no longer so plausible: it is a case of “extraordinary claims [requiring] extraordinary evidence” Sagan (1986).
Comparing scenarios item ii) and item iii) is also interesting. Overall, they suggest that it is still possible that a confounder is responsible for a large effect; however, this is plausible only if the confounder $Z$ can bias the estimate of the $T\to E$ effect with only a small change in the treatment $T$ (scenario item ii), where the $T\to E$ is more sensitive), as opposed to a large change (which is less likely to have been missed by the quantitative analysis).
As usual, with a better understanding of the domain (for example, the typical project characteristics in the company where the analyzed data was collected), the E-value analysis could produce more actionable results. For example, if one is confident that the estimates of team size and project effort that are normally produced are usually precise, it would indicate that it’s unlikely that a confounder would go undetected even for small changes of treatment/outcome (e.g., $\delta=0.1)$. Conversely, if the characteristics of the projects that are being analyzed make estimates intrinsically imprecise or uncertain, an unmeasured confounder that substantially increases the probability of swaying such estimates is a plausible possibility. As one example, Morasca and Russo (2001) study of productivity (one of those surveyed in Feldt et al. (2025)) presents estimates of productivity with a large dispersion ($\sigma\in[0.35,7.2]$); these indicate much uncertainty, thus raising the possibility of unmeasured confounders.
@@ 5 Dealing with Omitted Variable Bias
What kind of information we can use to build a DAG depends on the domain we are investigating, and on the maturity of the state of the art. If there are plenty of rigorous primary studies about the quantities of interest, and perhaps even systematic reviews or meta-studies, we can summarize their evidence in a DAG—similarly to what is proposed by Feldt et al. (2025). If our study’s target is novel or less established, we may have to rely on domain expertise and intuition to build a DAG. Even if there is a lot of uncertainty about the precise causal structure underlying a certain domain, there are techniques to (partially) validate candidate DAGs (Furia et al., 2023). It is quite natural to also consider different possible DAGs, and to use them to perform a “what if” analysis in different scenarios. In this step of the analysis, causal DAGs are mainly a convenient notation to rigorously express our knowledge or hypotheses about the causal relations among variables, and to investigate their consequences on the overall results of our analysis.
@@
As shown in subsection 4.2, given a DAG and treatment/outcome variables, one can systematically compute an adjustment set: a set of covariates that should be conditioned on in a regression to ensure that the coefficient associated with the treatment variable estimates the unbiased effect of the treatment on the outcome (“controlling for” the spurious influence of any confounders).
In the best-case scenarios, one can simply use an adjustment set to prevent confounding. Unfortunately, this is not always possible. First, the adjustment set’s validity is predicated on the accuracy and completeness of the DAG: if we missed some relevant variables, or misrepresented some causal relations, an adjustment set no longer guarantees an unbiased estimate. Second, even if we are confident the DAG is accurate, certain DAGs do not admit adjustment sets (because different kinds of confounding require incompatible adjustment sets (McElreath, 2020)), or some variables in the adjustment set are hard or impossible to measure (a scenario that we explored in subsection 4.3). In these cases, the next steps in this list can help deal with these shortcomings.
@@
Using the parameter estimates identified in the previous step, one can perform different kinds of sensitivity analyses of unmeasured confounder. These analyses provide a quantitative estimate (usually on a normalized scale) of how much some unmeasured confounder may bias the estimate of an effect of interest. In the paper, we demonstrated two kinds of so-called tipping-point sensitivity analysis, which express how strong an unmeasured confounder should be to cancel out an observed treatment/outcome effect. subsection 3.2 presented a tipping-point analysis based on an estimate of the scaled-mean difference of an unmeasured confounder on the treatment, which is applicable to dichotomous treatment variables. subsection 4.5 presented a sensitivity analysis based on E-values—a probability ratio between the confounded and non-confounded scenarios—which is also applicable to continuous treatment variables.
@@
If the previous step’s sensitivity analysis is inconclusive, in that it failed to rule out the possibility of confounding, one can perform a more precise analysis of confounding based on simulation. subsection 4.4 illustrated this on our second illustrative example of teamwork productivity. A simulation analysis is flexible, because one can explore many different variants of generative and inference models. It also supports analyzing the impact of dealing with small sample sizes in a way that realistically reflects the availability of data in the study domain. These advantages come with a cost in terms of simulation time; usually, however, it still is much cheaper to perform a detailed simulation than to embark “blind” in running an empirical study without a clear understanding of the possible confounders, and of the threats to validity they may introduce.
@@
5.1 Reporting a Sensitivity Analysis
To encourage researchers to perform sensitivity analyses as part of their quantitative studies, we outline a few simple guidelines on how to report the results of a sensitivity analysis in a paper.
DAG:
The DAG is a succinct summary of the key assumptions that underlie a sensitivity analysis. Therefore, it’s always recommended to include the DAG in the paper, preferably with a brief justification for its structure—or a reference to another publication that justifies it.
Parameter estimates:
The other key ingredient for a sensitivity analysis is an estimate of the main relations among variables—in particular, possible unmeasured confounders. Since there is usually a degree of uncertainty about the “real” parameter values, one should normally indicate a range of values for each parameter, rather than a single point estimate. Each estimate should also be accompanied by a succinct justification: a study that suggests that value, or, at least, a rationale for an educated guess. Table 4 shows an example of how the various parameter estimates could be reported in tabular form.
Tipping plot:
The outcome of a sensitivity analysis can be visually summarized with a plot of the tipping points similar to Figure 5 and 11(a). As we argued throughout the paper, a sensitivity analysis explores and compares different scenarios. Therefore, its outcome is generally not dichotomous but rather a matter of degree. A plot provides a concise, visual summary of “what if” scenarios such as: “if the unmeasured confounders are within certain ranges, then their confounding effect would/would not be able to nullify the observed effect”.
@@ 6 Conclusions
This paper introduced to the empirical software engineering community techniques to assess and mitigate the so-called omitted variable bias. These techniques are crucially based on a causal model of the relations among variables of interest, formalized by means the causal DAG notation (Pearl, 2009).
First, if the causal structure among variables admits an adjustment set that only includes measurable variables, one can correct for an omitted variable bias by simply conditioning on all variables in the adjustment set. If this is not possible, one can perform a sensitivity analysis, whose goal is investigating the impact of unknown confounders. The paper presented different kinds of sensitivity analyses, including both so-called tipping-point analyses based on canonical distributional assumptions, and more precise, but also computationally expensive, analyses based on simulation. We demonstrated these techniques on two illustrative examples—the relation between programming languages and code quality, and the effect on team size on software development effort—taken from recent statistical analyses of data in these two domains (Ray et al., 2014; Feldt et al., 2025).